# Rethinking Data Protection in the (Generative) Artificial Intelligence Era

### v1

@@ Abstract
The (generative) artificial intelligence (AI) era has profoundly reshaped the meaning and value of data. No longer confined to static content, data now permeates every stage of the AI lifecycle—from the training samples that shape model parameters to the prompts and outputs that drive real-world model deployment. This shift renders traditional notions of data protection insufficient, while the boundaries of what needs safeguarding remain poorly defined. Failing to safeguard data in AI systems can inflict societal and individual, underscoring the urgent need to clearly delineate the scope of and rigorously enforce data protection. In this perspective, we propose a four-level taxonomy, including non-usability, privacy preservation, traceability, and deletability, that captures the diverse protection needs arising in modern (generative) AI models and systems. Our framework offers a structured understanding of the trade-offs between data utility and control, spanning the entire AI pipeline, including training datasets, model weights, system prompts, and AI-generated content. We analyze representative technical approaches at each level and reveal regulatory blind spots that leave critical assets exposed. By offering a structured lens to align future AI technologies and governance with trustworthy data practices, we underscore the urgency of rethinking data protection for modern AI techniques and provide timely guidance for developers, researchers, and regulators alike.
@@ I Introduction
However, in the AI era, especially with the emergence of generative AI models, the scope of data protection has become far more complex and ambiguous [21, 22]. As shown in Figure 1, data permeates every stage of an AI model’s life cycle, making its value increasingly tied to the model rather than just the raw content of the data. For instance, developers compile many individual samples into large training datasets that feed into model development. The trained models themselves then become valuable data assets with significant commercial value. In addition, high-value or sensitive data ($e.g.$, original artworks or personal medical records) may also be incorporated as inputs during a model’s inference stage. Besides, with the rise of generative AI models, the outputs of inference are no longer simple predictions – they can be substantial content in their own right. For example, an LLM might generate executable code for a requested function [23], or a diffusion model might produce a realistic image for an advertisement or animation clips [24]. These AI-generated outputs are also valuable forms of data that merit protection.
@@
To tackle this problem, this paper offers the first timely overview and categorization of data protection in the (generative) AI era. Specifically, we introduce a hierarchical taxonomy of data protection comprising four distinct levels: data non-usability, privacy-preservation, traceability, and deletability. Each level in this taxonomy reflects a different balance between how usable the data remains for AI models and the degree of control or protection imposed on that data. At the most stringent end of the spectrum, data non-usability ensures that certain data cannot be used for model training or inference at all, offering maximal protection by completely sacrificing utility. Progressing down the hierarchy, privacy preservation permits data to be utilized in model development and application while safeguarding sensitive information, a trade-off that maintains some utility but enforces confidentiality of personal or private attributes. Further along, traceability allows nearly full data usage, yet embeds mechanisms to track the data’s origin and usage, thereby enabling transparency and accountability (for instance, detecting if data has been misused) with only minimal impact on the data’s functionality. Finally, at the most permissive level, data deletability lets data be fully integrated on the condition that its influence can be later removed from the model upon request. This last level emphasizes post-hoc control (aligning with ‘right to be forgotten’ principles) without impeding initial data utility. In particular, to ground this taxonomy, we systematically review representative technical approaches at each level, highlighting their strengths and limitations in practical settings.
@@ III Techniques for Data Protection
Techniques for Non-usability. Non-usability encompasses strategies that block any unauthorized party from using or even accessing protected data. Arguably, the most direct method is encryption [78, 79]: by securing data with strong cryptographic keys, the information remains unintelligible without proper authorization. A complementary line of defense ensures that the data cannot be exploited even if an adversary obtains it. For example, authorization mechanisms, including fine-grained data-access control [52, 80, 53] and model-level authorization [81, 82], allow only approved entities to obtain (correct) model outputs. Unauthorized requests receive degraded or nonsensical responses. Beyond controlling access or general utility, a further class of techniques makes the data unusable in unauthorized domains: unlearnable examples [32, 67] embed imperceptible perturbations that frustrate a model’s ability to extract ‘task-relevant’ features, whereas non-transferable learning [83, 84] deliberately suppresses ‘task-agnostic’ features so that any knowledge gleaned cannot generalize to unintended tasks. Together, these techniques align with a ‘secure-by-design’ philosophy: the data would be essentially non-usable and remain protected even in worst-case scenarios.
@@ IV Regulations on Data Protection
In the era of (generative) AI, regulation plays a foundational role in safeguarding data integrity, privacy, and accountability. Unlike traditional data systems, where protection focuses on static storage and access control, AI systems rely on dynamic, model-centric data use: once data is absorbed into a model’s parameters, it may persist, influence downstream outputs, and defy straightforward removal. Legal frameworks thus serve as critical instruments to constrain unauthorized use, enforce privacy preservation, ensure traceability, and empower users with redress and deletion rights. As shown in Table I, there are already some pioneering related regulations. For instance, many privacy laws operationalize L1 (Non-usability) by prohibiting the use of sensitive or unlawfully collected data for AI training altogether. Similarly, L2 (Privacy-preservation) is widely mandated through consent requirements, anonymization, and processing limits. Emerging regulations now also touch on L3 (Traceability)—requiring documentation of data provenance and logging of model operations—and even aspire to L4 (Deletability), allowing individuals to remove their data’s influence post-training. As the diffusion of data across AI pipelines complicates direct user control, the regulation remains the strongest binding force for aligning model development with ethical and societal norms.
@@ V Discussions
Data safety, in contrast, is primarily concerned with the content of the data and the potential harms arising from that content or the model’s behavior influenced by it [12, 131]. This includes issues like misinformation and deepfakes generated by models [132, 69], biases encoded in training data leading to discriminatory outcomes [133], the generation of harmful or illegal content, and the overall robustness and reliability of AI models and systems against adversarial manipulation aimed at causing malfunction or harm. In essence, data safety seeks to mitigate the negative consequences stemming from the data’s substance or the AI model’s outputs.

### v2

@@ Abstract
The (generative) artificial intelligence (AI) era has profoundly reshaped the meaning and value of data. No longer confined to static content, data now permeates every stage of the AI lifecycle—from the training samples that shape model parameters to the prompts and outputs that drive real-world model deployment. This shift renders traditional notions of data protection insufficient, while the boundaries of what needs safeguarding remain poorly defined. Failing to safeguard data in AI systems can inflict societal and individual harm, underscoring the urgent need to clearly delineate the scope of and rigorously enforce data protection. In this perspective, we propose a four-level taxonomy, including non-usability, privacy-preservation, traceability, and deletability, that captures the diverse protection needs arising in modern (generative) AI models and systems. Our framework offers a structured understanding of the trade-offs between data utility and control, spanning the entire AI pipeline, including training datasets, model weights, system prompts, and AI-generated content. We analyze representative technical approaches at each level and reveal regulatory blind spots that leave critical assets exposed. By offering a structured lens to align future AI technologies and governance with trustworthy data practices, we highlight the urgent need to rethink data protection for modern AI techniques and provide timely guidance for developers, researchers, and regulators alike.
@@ I Introduction
However, in the AI era, especially with the emergence of generative AI models, the scope of data protection has become far more complex and ambiguous [21, 22, 23]. As shown in Figure 1, data permeates every stage of an AI model’s life cycle, making its value increasingly tied to the model rather than just the raw content of the data. For instance, developers compile many individual samples into large training datasets that feed into model development. The trained models themselves then become valuable data assets with significant commercial value. In addition, high-value or sensitive data ($e.g.$, original artworks or personal medical records) may also be incorporated as inputs during a model’s inference stage. Besides, with the rise of generative AI models, the outputs of inference are no longer simple predictions – they can be substantial content in their own right. For example, an LLM might generate executable code for a requested function [24], or a diffusion model might produce a realistic image for an advertisement or animation clips [25]. These AI-generated outputs are also valuable forms of data that merit protection.
@@
To tackle this problem, this paper offers the first timely overview and categorization of data protection in the (generative) AI era. Specifically, we introduce a hierarchical taxonomy of data protection comprising four distinct levels: data non-usability, privacy-preservation, traceability, and deletability. Each level in this taxonomy reflects a different balance between how usable the data remains for AI models and the degree of control or protection imposed on that data. At the most stringent end of the spectrum, data non-usability ensures that certain data cannot be used for model training or inference at all, offering maximal protection by completely sacrificing utility. Progressing down the hierarchy, privacy-preservation permits data to be utilized in model development and application while safeguarding sensitive information, a trade-off that maintains some utility but enforces confidentiality of personal or private attributes. Further along, traceability allows nearly full data usage, yet embeds mechanisms to track the data’s origin and usage, thereby enabling transparency and accountability (for instance, detecting if data has been misused) with only minimal impact on the data’s functionality. Finally, at the most permissive level, data deletability lets data be fully integrated on the condition that its influence can be later removed from the model upon request. This last level emphasizes post-hoc control (aligning with ‘right to be forgotten’ principles) without impeding initial data utility. In particular, to ground this taxonomy, we systematically review representative technical approaches at each level, highlighting their strengths and limitations in practical settings.
@@ III Techniques for Data Protection
Techniques for Non-usability. Non-usability encompasses strategies that block any unauthorized party from using or even accessing protected data. Arguably, the most direct method is encryption [78, 79, 80]: by securing data with strong cryptographic keys, the information remains unintelligible without proper authorization. A complementary line of defense ensures that the data cannot be exploited even if an adversary obtains it. For example, authorization mechanisms, including fine-grained data-access control [53, 81, 54] and model-level authorization [82, 83], allow only approved entities to obtain (correct) model outputs. Unauthorized requests receive degraded or nonsensical responses. Beyond controlling access or general utility, a further class of techniques makes the data unusable in unauthorized domains: unlearnable examples [33, 68] embed imperceptible perturbations that frustrate a model’s ability to extract ‘task-relevant’ features, whereas non-transferable learning [84, 85] deliberately suppresses ‘task-agnostic’ features so that any knowledge gleaned cannot generalize to unintended tasks. Together, these techniques align with a ‘secure-by-design’ philosophy: the data would be essentially non-usable and remain protected even in worst-case scenarios.
@@ IV Regulations on Data Protection
In the era of (generative) AI, regulation plays a foundational role in safeguarding data integrity, privacy, and accountability. Unlike traditional data systems, where protection focuses on static storage and access control, AI systems rely on dynamic, model-centric data use: once data is absorbed into a model’s parameters, it may persist, influence downstream outputs, and defy straightforward removal. Legal frameworks thus serve as critical instruments to constrain unauthorized use, enforce privacy-preservation, ensure traceability, and empower users with redress and deletion rights. As shown in Table I, there are already some pioneering related regulations. For instance, many privacy laws operationalize L1 (Non-usability) by prohibiting the use of sensitive or unlawfully collected data for AI training altogether. Similarly, L2 (Privacy-preservation) is widely mandated through consent requirements, anonymization, and processing limits. Emerging regulations now also touch on L3 (Traceability)—requiring documentation of data provenance and logging of model operations—and even aspire to L4 (Deletability), allowing individuals to remove their data’s influence post-training. As the diffusion of data across AI pipelines complicates direct user control, the regulation remains the strongest binding force for aligning model development with ethical and societal norms.
@@ V Discussions
Data safety, in contrast, is primarily concerned with the content of the data and the potential harms arising from that content or the model’s behavior influenced by it [12, 131]. This includes issues like misinformation and deepfakes generated by models [132, 70], biases encoded in training data leading to discriminatory outcomes [133], the generation of harmful or incorrect hazardous content [134], and the overall robustness and reliability of AI models and systems against adversarial manipulation aimed at causing malfunction or harm [135]. In essence, data safety seeks to mitigate the negative consequences stemming from the data’s substance or the AI model’s outputs.